how to solve y''-y'-12y=e4x

1
Expert's answer
2021-06-16T11:31:26-0400
"y''-y'-12y=e^{4x}"

Write the related homogeneous or complementary equation:


"y''-y'-12y=0"


The general solution of a nonhomogeneous equation is the sum of the general solution "y_h(x)" of the related homogeneous equation and a particular solution "y_p(x)" of the nonhomogeneous equation:


"y(x)=y_h(x)+y_p(x)"

Consider a homogeneous equation 


"y''-y'-12y=0"

Write the characteristic (auxiliary) equation:


"r^2-r-12=0"

"(r-4)(r+3)=0"

"r_1=4, r_2=-3"

The general solution of the homogeneous equation is


"y_h(x)=C_1e^{4x}+C_2e^{-3x}"



Method of undetermined coefficients

Let the general solution of a second order homogeneous differential equation be


"y(x)=C_1(x)e^{4x}+C_2(x)e^{-3x}"

The unknown functions "C_1(x)" and "C_2(x)" can be determined from the system of two equations:


"C_1'e^{4x}+C_2'e^{-3x}=0"

"C_1'(4e^{4x})+C_2'(-3e^{-3x})=e^{4x}"



"C_2'e^{-3x}=-C_1'e^{4x}"

"4C_1'e^{4x}+3C_1'e^{4x}=e^{4x}"


"C_1'=\dfrac{1}{7}"

"C_2'=-\dfrac{1}{7}e^{7x}"





"C_1=\dfrac{1}{7}x+c_3"

"C_2=-\dfrac{1}{49}e^{7x}+c_2"



The general solution of a second order homogeneous differential equation be


"y(x)=c_1e^{4x}+c_2e^{-3x}+\dfrac{1}{7}xe^{4x}"